OpenStudy (julie):

Calculate the mole fraction of HCL, MM = 36.46 g/mol , in a 1.20 M aqueous, MM H20 = 18.02 g/mol , solution when a density is 1.18 g/mol 25 degree C.

OpenStudy (trump):

You know that 1L of the solution contains 1.20 moles HCl, to find the mole fraction, you must find how many moles of water the solution contains, and then divide the number of moles HCl by the total number of moles present in the solution. Here you have that 1 L of solution contains 1.20 mol HCl and it weighs 1.18g*1000mL=1180 g. The weight of water is 1180g-(36,46g/mol*1.20 mol HCl)=1180 g-43.75g=1136.25g water. Then divide the weight of the water by the Mm of H2O to find the number of moles H2O: 1136/18.02=63.05 mol. The last step is to divide the number of moles HCl by the total number of moles present in the solution: 1.20/(1.20+63.05)=0.1867, which is approximately 0.0187.